# Fragile Preferences: A Deep Dive into Order Effects in Large Language Models

### v1

@@ Abstract
Large language models (LLMs) are increasingly used in decision-support systems across high-stakes domains such as hiring and university admissions, where decisions often involve selecting among competing alternatives. While prior work has noted positional order biases in LLM-driven comparisons, these biases have not been systematically dissected or linked to underlying preference structures. We provide the first comprehensive investigation of positional biases across multiple LLM architectures and domains, uncovering strong and consistent order effects—including a novel centrality bias not previously documented in human or machine decision-making. We also find a quality-dependent shift: when options are high quality, models exhibit primacy bias, but favor latter options when option quality is low. We further identify a previously undocumented bias favoring certain names over others. To distinguish superficial tie-breaking from true distortions of judgment, we introduce a framework that classifies pairwise preferences as robust, fragile, or indifferent. We show that order effects can lead models to select strictly inferior options, and that positional biases are typically stronger than gender biases. These findings suggest that LLMs are not merely inheriting human-like biases, but exhibit distinct failure modes not seen in human decision-making. We propose targeted mitigation strategies, including a novel use of the temperature parameter, to reduce order-driven distortions.
@@ 1 Introduction
Large language models (LLMs) are increasingly integrated into decision-support systems across high-stakes domains such as hiring, healthcare, and loan approvals [1, 2, 3]. In these contexts, ensuring fairness and transparency is not just an ethical imperative but often a legal requirement, as recent regulations emphasize algorithmic accountability and nondiscrimination [4, 5, 6]. Among these domains, hiring has received particular attention due to both its societal importance and the real-world use of LLMs in resume screening [7, 6].
@@
To date, most studies of LLMs in hiring have focused on scoring individual resumes or responses, documenting biases related to gender, race, and socioeconomic background [8, 7, 9, 10, 11, 12, 13]. A growing line of work now advocates for pairwise or setwise comparisons, arguing that relative judgments are more consistent and more human-like [14, 15, 16, 17]. However, as these frameworks gain traction, it becomes increasingly urgent to examine the biases they may introduce, especially in high-stakes settings where even small distortions can have large consequences.
@@
Ranking-based decisions are particularly susceptible to positional and contextual effects. In human judgment, these include primacy and recency [18, 19, 20, 21, 22, 23], anchoring [24, 25], contrast effects [26, 27], and decoy effects [28, 29, 30]. These biases can undermine consistency and lead to inconsistent or irrational preferences. Recent studies show that LLMs are vulnerable to similar biases. Positional effects have been documented in multiple-choice settings [31, 32], classification tasks [33, 34], and pairwise evaluations [17, 35]. Other context-related effects, such as anchoring and decoy effects, have also been observed [36, 37]. Yet this literature remains largely descriptive: we know these effects exist, but not how they manifest in different contexts, how they interact with other biases, or what mechanisms drive them.
@@
In this work, we systematically investigate positional biases in LLMs. We uncover previously unreported biases, show that these effects can meaningfully distort underlying preferences, and find important differences from human biases. Across two domains—hiring and color selection—we uncover a consistent, quality-dependent pattern in pairwise comparisons: when the options are high quality, LLMs exhibit a primacy bias, favoring the first candidate or color. For lower-quality options, however, they favor latter options. In triplewise comparisons, we identify a novel centrality bias: a consistent preference for the middle option. Both the centrality effect and the quality-dependent shift from primacy to recency appear to be unique to LLMs and have not, to our knowledge, been reported in human (or LLM) decision-making. We further show that positional biases are, for the most part, stronger than gender biases, and uncover a strong and previously undocumented source of distortion: a bias favoring certain names over others. These patterns suggest that LLMs are not merely inheriting human-like heuristics from their training data, but are manifesting new failure modes with distinct underlying mechanisms.
@@
Separate from these novel patterns, we highlight a conceptual distinction that has been largely overlooked in prior work on order effects in both humans and LLMs. Positional biases are typically treated as a single phenomenon, but they can arise from two fundamentally different mechanisms: tie-breaking heuristics and true preference distortions. Tie-breaking occurs when the agent (human or AI) has no clear preference and defaults to a fixed positional rule (e.g., always guessing option C on a multiple-choice test). In contrast, a preference distortion occurs when the order of presentation alters an otherwise strict preference. This distinction matters: while tie-breaking is relatively harmless, preference distortion implies the model may select inferior options due solely to presentation order.
@@
Disentangling these mechanisms is difficult, especially in domains where ground-truth preferences are ambiguous. Many prior studies suggest that order effects emerge mainly when options are similar [31, 17, 32], consistent with the tie-breaking view. However, we show that positional effects in LLMs can lead to genuine preference reversals. Leveraging the ability to repeatedly query LLMs and control their sampling behavior via temperature, we introduce a simple yet effective mitigation strategy: querying the model several times at higher temperature settings. When the temperature parameter is increased, the models’ output becomes more stochastic, and repeated sampling shows more variation. We can quantify the model’s preferences by the frequency with which the options are selected: when the probability of selecting option $a$ over $b$ at temperature 1 is statistically significant, we infer that the model prefers $a$ to $b$, even if the model appears indifferent at temperature 0. In such scenarios, increasing the temperature can improve accuracy, contrary to the prevailing understanding that greater accuracy is attained at lower temperature settings[38, 39]. We elaborate on this technique in Subsection 2.3.
@@
To formalize the distinction between tie-breaking and true distortions, we extend the classical framework of decision-making, which represents preferences using the binary relations $a\succ b$ (strict preference) and $a\sim b$ (indifference). We introduce a refined notation that distinguishes between two types of strict preferences. We write $a\ggcurly b$ to indicate that the preference for $a$ over $b$ is strong and stable across contexts. In contrast, $a\succcurlyeq b$ denotes a genuine but unstable preference that may be reversed by order effects. This framework allows for clearer interpretation of observed behavior. if $a\succcurlyeq b$, but $b$ is consistently selected when presented first, we attribute the reversal to a distortion of the underlying preference, not to heuristic tie-breaking. However, if $a\sim b$, the cause for the order effect cannot be disambiguated.
@@
Together, our findings suggest that LLM comparisons are governed by both familiar and novel biases, and that these biases can systematically skew decisions. The implications are broad: positional distortions often outweigh demographic biases, and they cannot be reliably corrected by averaging or randomizing presentation. In our discussion, we explore both theoretical explanations and practical mitigation strategies.
@@ 2 Results
(a) Proportion of color choices by position across 4 quality tiers in pairwise comparisons. Tier quality is decreasing from left to right: ‘Plain’ is the highest quality tier and ‘Harsh’ is the lowest.
(b) Positional bias in pairwise resume selection, aggregated across professions.
Figure 1: Positional bias in pairwise comparisons at $T=1$. (a) shows the effect of presentation order on color selection in pairwise comparisons for each quality tier. (b) shows aggregate positional effects in resume evaluations across professions. In both tasks, higher-quality options tend to exhibit a primacy effect and lower-quality ones a recency effect, though the precise threshold separating them varies by model and domain. Full results for color comparisons are provided in Appendix C ; results for resumes disaggregated by profession are given in Appendix D
We evaluated positional biases in three widely used large language models—GPT-4o-mini, Claude 3 Haiku, and Llama 3 8B—across two domains designed to reflect real-world comparative decisions: resume screening and color selection. As order effects are most likely to arise when alternatives are similar in quality [17, 31], we constructed sets of items that are close in overall quality and grouped them into four tiers: ‘Ideal’, ‘Fair’, ‘Plain’ and ‘Harsh’ for colors and ‘Best’, ‘Good’, ‘Mediocre’ and ‘Weak’ for resumes. In the hiring domain, we used LLM-generated resumes spanning the four quality tiers in four professions. In the color domain, we curated sets of wall paint colors grouped by their suitability for a child’s room.
For each model and domain, we conducted exhaustive pairwise and triplewise comparisons, testing all permutations of item order. Prompts included “Select the strongest candidate” and “Which color is best for a kid’s room?” Each model was evaluated at multiple temperature settings. At temperature 0 ($T=0$), models behave almost deterministically, [^1] While temperature 0 is often assumed to produce deterministic outputs, we observed occasional variability in all models even at $T=0$. selecting the most likely answer. At temperature 1 ($T=1$), responses vary across repeated queries, enabling statistical analysis of positional effects. In our analyses, we treated each $T=1$ response as an independent sample from the model’s output distribution. Full experimental details appear in Section 4 (Materials and Methods) and the Appendix.
@@ 2.1 Order Effects in Pairwise Comparisons
We queried GPT-4o-mini, Claude 3 Haiku, and Llama 3 8B on exhaustive pairwise comparisons in two domains—resumes and colors—at multiple temperature settings. In every setting, higher-quality tiers exhibited a primacy effect (more likely to be selected when presented first), while lower-quality tiers showed a recency effect (more likely to be selected when presented second); these effects are clearly visible in Figure 1. For example, when comparing the two colors of the ‘Fair’ tier (Gentle Coral and Buttercream Yellow), GPT-4o-mini chose the first option 100% of the time at $T=0$ and 89% at $T=1$ (binomial $p=2.5\times 10^{-16}$). The quality tier at which bias flipped from primacy to recency was highest for Claude 3 Haiku and lowest for Llama 3 8B, a pattern consistent across both domains. Full statistics appear in Appendices C (colors) and D (resumes).
@@
(b) Aggregate positional bias in triplewise resume selection across 4 quality tiers.
Figure 2: Positional bias in triplewise comparisons at $T=0$. (a) shows the effect of presentation order on color selection. (b) shows aggregate positional effects in resume evaluations across professions. Additional results for color selection are provided in Appendix F, and resumes selection results disaggregated by profession are in Appendix G.
@@ 2.2 Order Effects in Triplewise Comparisons
In triplewise comparisons, higher-quality tiers exhibit a primacy effect, consistent with the pairwise results. For lower-quality tiers, all models favor latter positions, though they differ in whether the second or third option is preferred; see Figure 2. GPT-4o-mini and Claude 3 Haiku both display a centrality bias—favoring the middle option over first or last. For example, at $T=0$, GPT-4o-mini always selected the middle option for ‘Plain’ colors, and Claude 3 Haiku did so 83% of the time for ‘Harsh’ colors. In contrast, Llama 3 8B did not display a centrality bias, but instead showed simultaneous primacy and recency effects, resembling the serial position effects observed in humans [21, 40]: among colors in the ‘Fair’ tier at $T=0$, Llama 3 8B chose the first and last positions 50% of the time each, and never selected the middle option. See Appendices F (colors) and G (resumes) for additional details.
Across models and domains, the bias generally shifts monotonically toward later positions as quality declines, with the exception of Claude 3 Haiku in the color domain. Positional biases are notably weaker in resume selection for Claude 3 Haiku and Llama 3 8B compared to color selection. As we show in the Name Bias section below, this attenuation is largely due to interference from strong preferences for specific names.
@@
[…]
(a) Name selection frequencies in triplewise comparisons. The histogram shows how often each candidate name was selected across all triplewise tasks: selection counts reflect the number of times the name was chosen as the best option. See Appendix H for additional results and details.
(b) Gender and order effects at $T=0$ in pairwise resume comparisons. All models show stronger biases for presentation order than for gender. Results by profession and temperature are provided in Appendix I .
@@
[…]
2.4 Interaction of Positional Bias with Other Biases
In addition to positional order effects, LLMs exhibit other well-documented biases, such as preferences based on gender or identity. In this section, we examine how these biases interact with positional effects. While gender bias in algorithmic decision-making is well documented, we also uncover an additional bias: consistent preferences for specific individual names, even when those names are carefully selected to avoid known bias-triggering cues such as race and ethnicity. To our knowledge, such name-specific biases have not been documented in human decision-making under comparable conditions. Across all models, we find that these biases—though model-specific in strength and direction—are generally weaker than the positional effects, although in some cases they appear to compete with or modulate them.

### v2

@@ Abstract
Large language models (LLMs) are increasingly deployed in decision-support systems for high-stakes domains such as hiring and university admissions, where choices often involve selecting among competing alternatives. While prior work has noted position order biases in LLM-driven comparisons, these biases have not been systematically analyzed or linked to underlying preference structures. We present the first comprehensive study of position biases across multiple LLMs and two distinct domains: resume comparisons, representing a realistic high-stakes context, and color selection, which isolates position effects by removing confounding factors. We find strong and consistent order effects, including a quality-dependent shift: when all options are high quality, models favor the first option, but when quality is lower, they favor later options. We also identify two previously undocumented biases in both human and machine decision-making: a centrality bias—favoring the middle position in triplewise comparisons—and a name bias, where certain names are favored despite controlling for demographic signals. To separate superficial tie-breaking from genuine distortions of judgment, we extend the rational choice framework to classify pairwise preferences as robust, fragile, or indifferent. Using this framework, we show that order effects can lead models to select strictly inferior options, and that position biases are typically stronger than gender biases. These results indicate that LLMs exhibit distinct failure modes not documented in human decision-making. We also propose targeted mitigation strategies, including a novel use of the temperature parameter, to recover underlying preferences when order effects distort model behavior.
Keywords Large Language Models, Position Bias, Fragile Preferences, Preference Distortion
@@ 1 Introduction
Large language models (LLMs) are becoming increasingly integrated into high-stakes decision-support systems. For instance, in healthcare, LLMs are being used to assist in differential diagnosis and support triage decisions, where biases could have life-altering consequences [1]; in finance, LLMs support fraud detection and automated customer service, where biases could lead to significant economic harm [2]. In such contexts, ensuring fairness and transparency is not just an ethical imperative but often a legal requirement, as recent regulations emphasize algorithmic accountability and nondiscrimination [3, 4, 5]. Among these high-stakes domains, hiring stands out as a particularly relevant setting for studying AI biases: employment decisions have large economic and societal impacts, and LLMs are already being used extensively in multiple steps of the hiring process [6, 7, 5].
@@
To date, most studies of LLMs in hiring decisions have focused on documenting demographic biases in the scoring of individual resumes or responses [8, 7, 9, 10, 11, 12, 13]. A growing line of work now advocates for using pairwise or setwise comparisons in place of individual scoring, arguing that relative judgments are more consistent and human-like [14, 15, 16, 17]. As these frameworks become more widespread, it becomes increasingly urgent to examine the biases they may introduce, especially in high-stakes settings where even small distortions can have large consequences.
@@
Ranking-based decisions are particularly susceptible to positional and contextual effects. In human judgment, these include primacy and recency [18, 19, 20, 21, 22, 23], anchoring [24, 25], contrast effects [26, 27], and decoy effects [28, 29, 30]. These biases can undermine consistency and lead to irrational preferences. Recent studies show that LLMs are vulnerable to similar biases. Order effects have been documented in multiple-choice settings [31, 32], classification tasks [33, 34], and pairwise evaluations [17, 35]. Other context-related effects, such as anchoring and decoy effects, have also been observed [36, 37]. Yet this literature remains largely descriptive: we know position biases exist, but not how different contexts influence them, the extent to which they are model-specific, or how they interact with other biases. It is also unclear whether they reflect tie-breaking or true preference distortions, or how they differ from position biases in humans.
In this work, we systematically investigate position biases in LLMs. [^1] These should be distinguished from a different type of position bias, where a model favors certain input positions regardless of token content, due to the interaction between attention mechanisms and positional embeddings [38]. In contrast, we study biases where models favor options based on their position in a list of choices. We uncover previously unreported biases, show that these effects can meaningfully distort underlying preferences, and identify important differences from human biases. We study these effects across two fundamentally different domains: hiring decisions and children’s room color selection. The hiring scenario is a high-stakes application with a concrete link to existing literature on LLM biases. In contrast, color selection is a low-stakes, previously unexplored application that is minimalistic, neutral and easily replicable. Because it is largely free of other well-documented biases such as gender or race, it allows us to isolate positional effects more precisely than in complex domains. For each domain, we conduct experiments using four quality tiers. We uncover a consistent, quality-dependent pattern in pairwise comparisons across both domains: when both options are of relatively high quality, LLMs tend to favor the first candidate or color; when the options are of lower quality, they tend to favor the second. Although the trends are similar across models, what constitutes “high” and “low” quality may vary across models. In pairwise resume comparisons for instance, Llama 3 8B favors the first candidate in the top two tiers and the second in the bottom two, while GPT-4o-mini favors the first candidate in the top tier, and the second in the bottom three.
This pattern extends to triplewise comparisons: when all three options are high-quality, LLMs tend to prefer the first options, whereas with lower-quality options, they favor later ones. Additionally, we identify a novel centrality bias: a preference for the middle option in certain scenarios. Both the centrality effect and the quality-dependent shift appear to be unique to LLMs and have not, to our knowledge, been reported in human (or LLM) decision-making.
@@
We further show that position biases are, for the most part, stronger than gender biases, and uncover a significant and previously undocumented source of distortion: a bias favoring certain names over others. These patterns suggest that LLMs are not merely inheriting human-like heuristics from their training data, but are manifesting new failure modes with distinct underlying mechanisms.
@@
In addition to these novel patterns, we highlight a conceptual distinction that has been largely overlooked in prior work on order effects in both humans and LLMs. Although rarely distinguished in prior work, position biases can arise from two fundamentally different phenomena: tie-breaking heuristics and true preference distortions. Tie-breaking occurs when the agent (human or AI) has no clear preference and defaults to a fixed positional rule (e.g., always guessing option C on a multiple-choice test). In contrast, a distortion occurs when the order of presentation alters an otherwise strict preference. This distinction matters: while tie-breaking is relatively harmless, preference distortion implies the model may select inferior options due solely to presentation order.
@@
Disentangling these mechanisms is difficult, especially in domains where ground-truth preferences are ambiguous. Many prior studies suggest that order effects emerge mainly when options are similar [31, 17, 32], consistent with the tie-breaking view. However, we show that positional effects in LLMs can lead to genuine preference reversals. Leveraging the ability to repeatedly query LLMs and control their sampling behavior via temperature, we introduce a simple yet effective mitigation strategy: querying the model numerous times at higher temperature settings.
When the temperature parameter (denoted $T$) increases, the models’ output becomes more stochastic, and repeated sampling shows more variation. For example, $T=0$ produces (nearly) deterministic outputs, while higher values such as $T=1$ yield greater randomness. We can determine the model’s underlying preferences by examining the frequency with which options are selected across multiple trials. When options are presented multiple times in all permutations at $T=1$, we infer that the model truly prefers option $a$ over $b$ if the probability of selecting $a$ is statistically significantly greater than that of selecting $b$ across all presentations. This approach reveals that some apparent indifference at $T=0$ actually masks underlying preferences that become detectable at higher temperatures.
In such scenarios, increasing temperature can improve preference detection accuracy, contrary to the prevailing understanding that greater accuracy is attained at lower temperature settings [39, 40]. We elaborate on this technique in Section 2.3 (Results).
@@
To formalize the distinction between tie-breaking and true distortions, we extend the classical rational choice framework, which represents preferences using the binary relations $a\succ b$ (strict preference) and $a\sim b$ (indifference). We introduce a refined notation that distinguishes between two types of strict preferences. We write $a\ggcurly b$ to indicate that the preference for $a$ over $b$ is robust and stable across contexts. In contrast, $a\succcurlyeq b$ denotes a fragile preference: a genuine but unstable preference that may be reversed by order effects.
We define these formally within our setting using random utility theory [41], where one defines preferences using selection probabilities. Let $\Pr[a|\{a,b\},t]$ denote the overall probability of selecting $a$ when both permutations $(a,b)$ and $(b,a)$ are presented an equal number of times at temperature $t$.
At $T=0$, the model’s output is (nearly) deterministic, so there are only two possible behaviors: (i) the same option is selected in both orders, giving selection probabilities of $\{0,1\}$, (ii) the model consistently picks one position (e.g., the first option), so each option is chosen half of the time, giving $\{0.5,0.5\}$. In practice, $T=0$ is not perfectly deterministic; we use $\approx$’ to indicate probabilities close to, but not exactly, 0, $0.5$, or $1$.
  • If $\Pr[a|\{a,b\},0]\!\approx\!1$ and $\Pr[a|\{a,b\},1]\!>\!0.5$ then $a\ggcurly b$.
  • If $\Pr[a|\{a,b\},0]\!\approx\!0.5$ and $\Pr[a|\{a,b\},1]\!>\!0.5$ then $a\succcurlyeq b$.
  • If $\Pr[a|\{a,b\},0]\!\approx\!0.5$ and $\Pr[a|\{a,b\},1]\!=\!0.5$ then $a\sim b$.
Preferences are deemed inconsistent if none of the above conditions hold; we have not observed any such cases. This framework allows for clearer interpretation of observed behavior: if $a\succcurlyeq b$, then order-dependent reversals (e.g., choosing $b$ when it appears first at $T=0$ despite an underlying preference for $a$) represent genuine preference distortions rather than tie-breaking heuristics. However, if $a\sim b$, the cause for the order effect cannot be disambiguated. Using this framework, repeated sampling at $T=1$ enables us to empirically distinguish between fragile preference and indifference, although in practice it may be difficult to separate them if $\Pr[a\mid\{a,b\},1]$ is very close to $0.5$.
@@
Our findings suggest that LLM comparisons are governed by both familiar and novel biases, and that these biases can systematically skew decisions. The implications are broad: positional distortions often outweigh demographic biases, and they cannot be reliably corrected by averaging or randomizing presentation. In our discussion, we explore both theoretical explanations and practical mitigation strategies.
@@ 2 Results
(a) Proportion of color choices by position across 4 quality tiers in pairwise comparisons. Tier quality is decreasing from left to right: ‘Ideal’ is the highest quality tier and ‘Harsh’ is the lowest.
(b) Position bias in pairwise resume selection, aggregated across professions.
[…]
We evaluated position biases in three widely used large language models—GPT-4o-mini, Claude 3 Haiku, and Llama 3 8B—across two domains selected to balance realism and control over confounding factors: resume screening and color selection. As order effects are most likely when alternatives are similar in quality [17, 31], we created item sets close in quality and grouped them into four tiers: Best, Good, Mediocre, and Weak for resumes, and Ideal, Fair, Plain, and Harsh for colors. In the hiring domain, we used LLM-generated resumes for four professions (Mechanical Engineer, Registered Nurse, Journalist, and Real Estate Agent). In the color domain, we curated sets of wall paint colors grouped by suitability for a child’s room.
For each model and domain, we conducted exhaustive pairwise and triplewise comparisons, testing all permutations of item order. Prompts included “Select the strongest candidate” and “Which color is best for a kid’s room?” Each model was evaluated at multiple temperature settings. At $T=0$, models behaved with near-determinism, [^2] While $T=0$ is often assumed to produce deterministic outputs, we observed occasional variability in all models. selecting the most likely answer. At $T=1$, responses varied across repeated queries, allowing statistical analysis of positional effects; each $T=1$ response was treated as an independent sample from the model’s output distribution.
We define positional biases as follows. A model exhibits a primacy bias if there exists a set of options such that the model disproportionately prefers whichever option is presented first. It exhibits a recency bias if there exists a set of options such that it disproportionately prefers whichever option is presented last. In triplewise comparisons, it exhibits a centrality bias if there exists a set of three options such that it disproportionately prefers whichever option is presented second. A model may exhibit more than one bias depending on the context. Full experimental details appear in Section 4 (Materials and Methods) and the Appendix.
For each model and domain, we conducted exhaustive pairwise and triplewise comparisons, testing all permutations of item order. Prompts included “Select the strongest candidate” and “Which color is best for a kid’s room?” Each model was evaluated at multiple temperature settings. At $T=0$, models behaved with near-determinism, selecting the most likely answer. At $T=1$, responses varied across repeated queries, allowing statistical analysis of positional effects; each $T=1$ response was treated as an independent sample from the model’s output distribution.
We define positional biases as follows: a model exhibits a primacy bias if it disproportionately prefers whichever option is presented first, a recency bias if it disproportionately prefers the last option, and a centrality bias (in triplewise comparisons) if it disproportionately prefers the middle option. A model may exhibit more than one bias depending on the context.
Full experimental details appear in Section 4 (Materials and Methods) and the Appendix.
@@ 2.1 Order Effects in Pairwise Comparisons
We queried the three LLMs on exhaustive pairwise comparisons in both domains at multiple temperature settings. In every setting, higher-quality tier(s) exhibited a primacy effect (more likely to be selected when presented first), while lower-quality tier(s) showed a recency effect (more likely to be selected when presented second). For instance, in color selection (1(a)), GPT-4o-mini displayed a primacy effect in the top two tiers (Ideal and Fair) and a recency effect in the bottom two (Plain and Harsh). Claude 3 Haiku and Llama 3 8B showed a primacy effect in the top one and three tiers, respectively, and a recency effect in the bottom three and one tiers, respectively. In 1(b), the resume selection results followed a similar overall pattern (though GPT-4o-mini and Claude 3 Haiku showed primacy only in the top tier, and Llama 3 8B in the top two tiers).
Full statistics, including those for additional temperature settings, appear in Appendices C (colors) and D (resumes).Consistent with our definitions of robust and fragile preferences, the results at $T=0$ are much more pronounced.
@@
(b) Aggregate position bias in triplewise resume selection across 4 quality tiers.
[…]
@@ 2.2 Order Effects in Triplewise Comparisons
In triplewise comparisons, higher-quality tier(s) exhibited a primacy effect, consistent with the pairwise results. For lower-quality tier(s), all models favored later positions, though they differed in whether the second or third option was preferred. Specifically, GPT-4o-mini and Claude 3 Haiku sometimes favored the second (middle) position, exhibiting centrality bias. For example, at $T=0$ (2(a)), GPT-4o-mini always selected the middle option for Plain colors, and Claude 3 Haiku did so 83% of the time for Harsh colors. In resume selection (2(b)), GPT-4o-mini preferred the middle option in three out of the four tiers tested.
In contrast, Llama 3 8B did not display a centrality bias but instead showed simultaneous primacy and recency effects, resembling the serial position effects observed in humans [21, 42]. Among colors in the Fair tier at $T=0$, Llama 3 8B chose the first and last positions equally often (50% each) and never selected the middle option. See Appendices F (colors) and G (resumes) for additional details.
Across models and domains, position bias generally shifted toward later positions as quality declined, with the exception of Claude 3 Haiku in the color domain. Position biases were notably weaker in resume selection for Claude 3 Haiku and Llama 3 8B compared to color selection. As we show in the Name Bias section below, this attenuation is largely due to interference from strong preferences for specific names.
To explore whether position biases arise in larger choice sets, we conducted preliminary fourwise color comparisons for each model. Although we evaluated only a single quality tier per model, the results demonstrate that position biases can persist in more complex choice scenarios. For full details, see Appendix H.
@@
We therefore repeated all comparisons at $T=1$, where increased stochasticity can reveal suppressed preferences. To build intuition, consider a stylized example with von Neumann–Morgenstern (VNM) utilities. [^3] A VNM utility function is any function that represents a preference relation consistent with the von Neumann–Morgenstern axioms. See [43] for a formal definition. We use VNM utilities only for intuition; we do not claim that the models’ preferences can be mapped to such a function or that they satisfy the axioms of rationality. Suppose the model assigns $u(a)=5$, $u(b)=5.1$, and there is a primacy boost $\delta=0.2$. At $T=0$, presenting $(a,b)$ yields $a$ (since $5+0.2>5.1$), and presenting $(b,a)$ yields $b$ (since $5.1+0.2>5$). Now suppose $T=1$ adds uniform noise $\varepsilon\sim U[-0.4,0.4]$ to the first option’s utility. [^4] Noise can be assumed to be added to only one position without loss of generality. Then $(a,b)$ yields $a$ with probability 0.625 ($\mathbb{P}[5.2+\varepsilon\geq 5.1]=0.625$), and $(b,a)$ yields $b$ with probability 0.875. Thus, $b$ is significantly more likely to be chosen than $a$ at $T=1$, revealing a strict (but fragile) preference for $b$ over $a$, rather than indifference.
Consistent with this intuition, eight of twelve model–tier color pairs in our experiments showed a significant preference ($p<0.05$), and three of these showed overwhelming distortion ($p<10^{-6}$). We obtained similar results for resumes; in one case, one resume was 1.5 times more likely to be chosen than the other, indicating that order effects can lead to the consistent selection of inferior options—effects that are most pronounced at lower temperatures. For more details, see Appendix E.
(a) Name selection frequencies in triplewise comparisons. The histogram shows how often each candidate name was selected across all triplewise tasks. The x-axis represents the number of times a name was selected, and the y-axis shows how many names were selected that number of times. For ChatGPT-4o-mini, nine names were selected between 30 and 32 times, and all names were selected between 24 and 38 times. See Appendix I for additional results and details.
(b) Gender and order effects at $T=0$ in pairwise resume comparisons. All models show stronger biases for presentation order than for gender. Results by profession and temperature are provided in Appendix J .
@@
2.4 Interaction of Position Bias with Other Biases
In addition to positional order effects, LLMs exhibit other well-documented biases, such as preferences based on gender or identity. In this section, we examine how these biases interact with positional effects. While gender bias in algorithmic decision-making is well documented, we also uncovered an additional bias: consistent preferences for specific individual names, even when those names were carefully selected to avoid known bias-triggering cues such as race and ethnicity. To our knowledge, such name-specific biases have not been documented in human decision-making under comparable conditions. Across all models, we found that both gender and name biases—though model-specific in strength and direction—are generally weaker than the positional effects, although in some cases they appear to compete with or modulate them.